Circular No. 8331

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2004ax, 2004aw, 2004bh, 2004bi, 2004bj, AND 2004bk

C. L. Gerardy, University of Texas at Austin; and R. A. Fesen, Dartmouth College, report that a low-resolution optical spectrogram (resolution 600; range 565–950 nm) of SN 2004bk (IAUC 8329), obtained on Apr. 24.28 UT with the 2.4-m Hiltner Telescope (+ ModSpec spectrograph) at MDM Observatory, shows it to be a type-Ia supernova near maximum light; adopting the 6906 km/s redshift of NGC 5246 from the NASA/IPAC Extragalactic Database (NED), the expansion velocity of the Si II feature (rest wavelength 635.5 nm) is 11800 km/s. A similar spectrogram of SN 2004bh (IAUC 8320), obtained on Apr. 24.21, shows it to be a type-II supernova, likely at early epoch. The spectrum exhibits a single broad Hα emission feature with little or no blueshifted absorption component. Adopting the redshift of UGC 5161 (10079 km/s) from the NED, the expansion velocity of the blue edge of the Hα feature is ∼ 6850 km/s. No other emission or absorption features are seen in the spectrum.

A. V. Filippenko, L. Desroches, M. Ganeshalingam, and R. Chornock, University of California, Berkeley; and F. J. D. Serduke, Lawrence Livermore National Laboratory, report that inspection of CCD spectra (range 330–1000 nm), obtained on Apr. 24 UT with the Shane 3-m telescope at Lick Observatory, reveals that SN 2004aw (IAUC 8310, 8311) is of type Ic, similar to SN 1991A (Filippenko 1992, Ap.J. 384, L37), rather than of type Ia as stated in IAUC 8312. SN 2004ax (IAUC 8311) is indeed a supernova, possibly of the ‘calcium-rich’ type-Ib/Ic variety like SN 2001co and SN 2003H (IAUC 8159). The Ca II near-infrared triplet emission is the only clear feature, although the data are noisy; a higher-quality spectrum should be obtained to verify the classification. SN 2004bh is of type II, perhaps II-L, at an early stage of development. The spectrum is dominated by Hα emission, with little corresponding blueshifted absorption. Weak Ca II near-infrared triplet absorption is also visible. SN 2004bi (IAUC 8321) appears to be of type IIb, not Ic as stated on IAUC 8324; the spectrum is similar to that of SN 1996cb (Qiu et al. 1996, A.J. 117, 736). SN 2004bj (IAUC 8329) is a type-Ia supernova, with a spectral-feature age (Riess et al. 1997 A.J. 114, 722) of 2 ± 2 days before maximum light. SN 2004bk is likewise of type Ia, but with a spectral-feature age of 5 ± 2 days after maximum light.

COMET C/2004 F4 (BRADFIELD) Corrigendum. On IAUC 8329, line 3, for p.a. 20^{o} read p.a. 290^{o}

2004 April 24  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green